OpenStudy (anonymous):

solve the system of equations by the substitution method 6x-3y=3 x+4y=0

OpenStudy (tyteen4a03):

Hint: use the variable that has no coefficient. It makes substitution easier. In this case, turn x + 4y = 0 into x = -4y. Then substitute x = -4y into 6x - 3y = 3.

OpenStudy (anonymous):

still lost please help

OpenStudy (tyteen4a03):

After substitution, the first equation becomes 6(-4y) - 3y = 3. Solve this equation and use the newfound value of y to solve for x.

OpenStudy (anonymous):

from 2nd equation get value of x and put in 1st equation and get value of x and put value of x in y.

OpenStudy (anonymous):

Solve for x in the second equation. It will be x=-4y. You get this by subtracting 4y to both sides of equation. Take x=-4y and plug it into 6x-3y=3. It will look like this 6(-4y)-3y=3, since x is -4y you can use that as the x in 6x. Distribute 6(-4y)-3y=3 which will be -24y-3y=3 or -27y=3. Solve for y which is y= 3/-27 = -9. So y=-9. Given that you found x, x=-4y, plug y=9 in. So x=-4(-9) which is x=36. Your answer might look like (36, -9) or x=36 and y=-9